VIBRATION-ROTATION-TUNNELING STATES OF THE BENZENE DIMER: AN AB INITIO STUDY

AD VAN DER AVOIRD, \underline{P. R. BUNKER}, MELANIE SCHNELL, GERT VON HELDEN, and GERARD MEIJER, Fritz-Haber-Institut der Max-Planck-Gesellschaft, Faradayweg 4-6, D-14195 Berlin, Germany; RAFAŁ PODESZWA and KRZYSZTOF SZALEWICZ, Department of Physics and Astronomy, University of Delaware, Newark, DE 19716; CLAUDE LEFORESTIER, Institut Charles Gerhardt Montpellier, Université Montpellier 2, 34095 Montpellier, Cedex 05, France; and ROB VAN HARREVELT, Theoretical Chemistry, Institute for Molecules and Materials, Radboud University Nijmegen, 6525 AJ Nijmegen, The Netherlands.

An improved intermolecular potential surface for the benzene dimer has been constructed from interaction energies computed by symmetry-adapted perturbation theory with the inclusion of third-order contributions. The potential was used in calculations of vibration-rotation-tunneling levels of the dimer by a method appropriate for large amplitude intermolecular motions and tunneling between multiple equivalent minima in the potential. The resulting levels were analyzed using the permutation-inversion full cluster tunneling group G_{576} and a chain of subgroups that starts from the molecular symmetry group C_{s}(M) of the rigid dimer at its equilibrium C_{s} geometry and leads to G_{576} if all possible intermolecular tunneling mechanisms are feasible.